Hi,

I've been working on this problem for a while, and am still getting an answer than doesn't work when I check/proof it. Here's how I've been trying to solve it:

5(3 - 2x) = 4(4x + 20) or
15 - 10x = 16x + 80 or
-65 = 26x or
x = -65/26

and then to proof
5(3 - 2(-65/26)) = 4(4-65/26)+ 20) or
5(78/26 - 130/26) = 4(260/26 + 520/26) or
-260/26 = 3120

The problem is:

"5(3 - 2x) = 4(4x + 20)"

It looks so simple... what am I doing wrong?

Thanks,
~ Jasmine

Answer by tutorcecilia(2152)

Your answer is correct:
5(3 - 2x) = 4(4x + 20) or
15 - 10x = 16x + 80 or
-65 = 26x or
x = -65/26
x=-2.5 [Decimal form is easier to check]
.
and then to proof
5(3 - 2(-2.5)) = 4(4(-2.5)+ 20) [Check by plugging -2.5 back into the equation]
40=40 [Checks out]